A bag contains six white, five red, and four balls. If six ball are drawn, find the probability that there will be two balls of each color.
Answer by stanbon(75887)

A bag contains six white, five red, and four balls. If six ball are drawn, find the probability that there will be two balls of each color.
---
# of ways to succeed:: 6C2 * 4C2 * 5C2 = 15*6*10 = 900
# of sets of 6 balls:: 15C6 = 5005
---
P(condition) = 900/5005 = 0.1798
-----------
Cheers,
Stan H.
---------------------